What is the pressure of 0.33 moles of nitrogen gas, if ita volume is 15.0 L at -25.0⁰C


1
Expert's answer
2021-05-31T02:06:33-0400

Using the ideal gas law "PV =nRT" , we find that the pressure will be "P =\frac{nRT}{V}" . Then, we'll substitute and find the pressure, using T = -25 °C = 248.15 K and R = 0.0821 "\frac{atm\cdot L}{mol \cdot K}" :


"P =\frac{nRT}{V} = \frac{(0.33\,\cancel{mol})(0.0821\frac{atm\cdot \cancel{L}}{\cancel{mol \cdot K}})(248.15\,\cancel{K})}{15.0\,\cancel{L}} = 0.4482\,atm"


In conclusion, the pressure of this gas is P=0.4482 atm.


Reference:

  • Chang, R. (2010). Chemistry. McGraw-Hill, New York.